the osmotic pressure of a solution is 3.2,which contains 45gm of sucrose in per liter of solution in 20 degree celcius calculate R
1
Expert's answer
2012-03-13T07:56:50-0400
PV=nRT
R=PV/nT
P=3.2
T=273+20=293
n=m/Mw=45/342=0,1315
v=1
l
R=3.2*1/293*0.1315=0.0830